FOURIER ANALYSIS AND PHYLOGENETIC TREES

STEVEN N. EVANS

Abstract. We give an overview of phylogenetic invariants: a technique for reconstructing evolutionary family trees from DNA sequence data. This method is useful in practice and is based on a number of simple ideas from elementary group theory, probability, linear algebra, and commutative algebra.